For what value of 'm' does the equation, {mx² + (3x-1)m + 2x +5=0} have equal roots but of opposite sign ??

Grade:10

2 Answers

Arun
4 years ago
Dear student
 
Use condition that f(0)
 
For x² + (3m+2)/m x + (5-m)/m = 0
 
Regards
Arun (askIITians forum expert)
Akash
4 years ago
If roots are equal in magnitude but opposite in sign, then sum of the roots = 0
Given equation can be written as x² + (3m+2)/m x + (5-m)/m = 0
Hence, sum of roots = (3m+2)/m = 0
Hence, m = – 2/3